Circular No. 8720

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006cs, 2006ct, 2006cu, 2006cv, 2006cw Five new apparent supernovae found on unfiltered CCD frames have

been reported: 2006cs by N. Y. Lee and W. Li (LOSS/KAIT; cf. IAUC 8716); 2006ct by R. Quimby and P. Mondol (cf. IAUC 8713); and the rest by the ‘Nearby Supernova Factory’ collaboration (cf. CBETs 538, 540).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006cs June 3.29 13^{h}45^{m}33.^{s}89 +35^{o}36′36′′.6 17.8 13′′.4 W, 3′′.8 S
2006ct May 25.21 12 09 57.03 +47 05 43.9 17.0 3′′.4 E, 1′′.4 S
2006cu May 29.4 14 47 43.31 + 9 39 33.9 18.1  –
2006cv June 5.4 14 47 47.83 +23 17 58.6 19.1  –
2006cw June 5.4 15 03 01.01 +21 41 40.6 19.4  –

Additional magnitudes for 2006cs in MCG +06-30-79: May 18.29 UT, [19.0; June 4.31, 17.7. A spectogram of 2006cs obtained on June 5.29 shows it to be a type-Ia supernova around maximum light (details on CBET 539). Additional magnitudes for 2006ct, which is also a type-Ia supernova (details on CBET 537): 2004 Dec. 15–2006 Jan. 28, [18.7 (co-addition of images); May 31.21, 16.9; June 3.19, 16.6. SNe 2006cu and 2006cv are type-IIn supernovae, and 2006cw is a type-II supernova (cf. CBET 538, 540).

COMETS C/2006 J4–J8 (SOHO) Additional comets have been found on SOHO website images (cf. IAUC

8719) — all being Kreutz sungrazers except C/2006 J5 (Meyer group). C/2006 J4 was stellar in C3 images, reaching mag 4.9 on May 8.762 UT at 9.1R_{⊙}; C2 images showed a very faint tail that reached a length of 500′′ on May 9.188 at 3.7R_{⊙}. C/2006 J5 was very small, stellar but slightly elongated in the direction of motion, and too faint for photometry. C/2006 J6, which was also found by J. Sachs, was extremely faint, diffuse, and elongated. C/2006 J7 and J8 were very close (only a few pixels apart), the leading component (J7) being marginally larger, brighter, and longerlasting; both components were small and stellar.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 J4 May 7.696 3^{h}16.^{m}6 +13^{o}35′ C3/2 KB 2006-L31
C/2006 J5  7.979 3 03.2 +18 16 C2  HS 2006-L31
C/2006 J6  8.854 3 07.3 +15 52 C2  RM 2006-L31
C/2006 J7  10.163 3 12.1 +16 13 C2  TH 2006-L32
C/2006 J8  10.163 3 12.1 +16 13 C2  TH 2006-L32
2006 June 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green